Operations Report November 10th

Crew 182 Operations Report 10Nov2017

SOL: 7

Name of person filing report: Carmen Atauconcha

Non-nominal systems:

WATER HEATER: It is not working. Today we wanted to take a shower and the water was cold. The heater is not heating the water. We opened the entire water pipe, but it still did not heat.

COMMUNICATION RADIOS: One of the radios is making a special noise constantly. Like every 3-5minutes it sounds, even if we leave it charging.
SECONDARY AIRLOCK: The door of the secondary airlock is gotten out-off-square.
ATV’s: The blue ATV is not working. When we try to turn it on, it does not answer.

Generator— Operating from 09 November, 2017 at 5pm and currently working. 26 hours.



8:00 AM – 17:00 PM 100%


8.43 am 98%
9.13 am 98%
9.43 am 98%
10.09 am 100%
10.39 am 100%
12.20 pm 100%
13.01 pm 98%
13.38 pm 96%
13.46 pm 95%
13.56 pm 94%
14.11 pm 93%
14.34 pm 100%
15.19 pm 100%
16.28 pm 98%
17.31 pm 93%

Diesel – 59 %

Propane – 51 psi

Gasoline (5 Gallon containers for ATV) – 1/4 gallons

Water (trailer) – 0 gallons

Water (static) – 275 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 66.8 gallons

Toilet tank emptied: yes


Oil Checked: yes

Atv # Fuel Used Gals: 1 gal

Tires Status: Perfect

# Hours the ATVs were Used today: 1 and ½ hours

Deimos used: Yes

Hours: 1 and a ½ hours

Beginning charge: 12:30 pm

Ending charge:

Currently charging: Yes

HabCar used? No

Notes and Comments:
Today we had an advance in the topic about recycling water. The number of gallons of water that we used today is less than the gallons of water used yesterday. And, I know that tomorrow is going to be better.
Summary of Hab operations:

We put the jumper cables delivered on the lower shelf in the EVA room. Also, we fixed the radios because some of them did not have a good sound. In addition, tomorrow we are going to receive new communication radios.

Summary of GreenHab operations:

We are going to heat the greenhouse tomorrow

Summary of ScienceDome operations:

We installed the autoclave perfectly. Even, it has already been used once without any difficulty. At the same time, tomorrow we are going to receive a new shop vac.

Summary of health and safety issues:

The HOS officer needs two fire alarms for the science dome and for the greenhouse.

Questions, concerns and requests to Mission Support:

We need you to repair the heater water, since it still does not work.

For the next week at the most on Monday, we are going to need water.

I have a doubt, which is the minimum limit for the propane tank and the diesel tank?

Where does the used water go?

Copyright © The Mars Society. All rights reserved. | Main Site